If potassium cyanide (KCN) reacts with acid, it will form a gas that is very toxic and can cause death, namely hydrogen cyanide gas (HCN). The chemical equation is: KCN (aq) + HCl (aq) → KCl (aq) + HCN (g) If a sample contains 0.140 g KCN is reacted with 100 mL of 0.1 M HCl. a. Determine what compound acts as the limiting reagent. (prove it by calculation). b. Calculate the mass of HCN produced in the above reaction
1
Expert's answer
2020-11-12T07:12:00-0500

Balanced reaction equation:

KCN + HCl = KCl + HCN

Quantity of KCN:

0.140 g / (65.12 g/mol) = 0.00215 mol

Quantity of HCl:

0.1 L * (0.1 mol/L) = 0.01 mol

Ratio of chemicals according to the equation is 1 : 1, so KCN is limiting reagent.

Quantity of HCN produced is 0.00215 mol

Mass of HCN:

0.00215 mol * (27.03 g/mol) = 0.058 g

Answer:

a) limiting reagent - KCN

b) mass of HCN - 0.058 g